gabrielleevans1:

Find the values of y = p(x) = √x for x = 0, 1.44, 2.25, 3.24, 4.41, 5.29

Liam:

p(0) = 0 p(1.44) = 6/5=1.2 p(2.25) = 3/2 = 1.5 p(3.24) = 9/5 =1.8 p(4.41) = 21/10 2.1 p(5.29) = 23/10=2.3